Global smartphone sales experienced extraordinary growth starting in 2007 with the introduction of both the iPhone and Android-platform based devices (see the Chapter Introduction). Following are the yearly global sales (in million units) for smartphones beginning in 2008 through 2015:

YEAR SALES

2008 140

2009 173

2010 305

2011 495

2012 730

2013 1030

2014 1300

2015 1430

Develop a linear trend line model to forecast smartphone sales for 2016, 2017, and 2018, and compare these forecasts with those reported on the Internet
Answer #1

Answer a: the trendline equation is y = 203.27x - 214.36

Explanation:

  • linear trend forecast- select the period and demand and insert a X-Y chart. Show trend line equation. To find forecast, replace the value of x in the equation with period no. and find the value of Y.

use the table below

Year Year no. X Sales, Y
2008 1 140
2009 2 173
2010 3 305
2011 4 495
2012 5 730
2013 6 1030
2014 7 1300
2015 8 1430

Answer b: For a forecast of year 2016, keep x= 9

F-2016 = y = 203.27*9 - 214.36 = 1615.07

For a forecast of year 2017, keep x= 10

F-2017 = y = 203.27*10 - 214.36 = 1818.34